Please help me to solve: 10-3x=2x-8x+40
Answer by rapaljer(4551)

Combine like terms on the right side:
10-3x=2x-8x+40
10-3x=-6x+40

Add +6x to both sides to get all x terms to the left side:
10 - 3x + 6x = -6x + 6x + 40
10 + 3x = 40

Subtract 10 from each side:
10 - 10 + 3x = 40 - 10
3x = 30
x= 10

R^2 at SCC